OpenStudy (anonymous):

Solve for x in x-3/2=x+1/5.

OpenStudy (anonymous):

\[\frac{ x-3 }{ 2 } = \frac{ x+1 }{ 5 }\]

OpenStudy (jhannybean):

Cross multiply.

OpenStudy (anonymous):

Would this be the equation? 5x-15=2x+2

OpenStudy (jhannybean):

Yes, now bring all your x's to one side of the equation.

OpenStudy (jhannybean):

So what would you get?

OpenStudy (anonymous):

17/3 which is 5.6?

OpenStudy (anonymous):

dont like mathhh

OpenStudy (anonymous):

i suck at math so dont make fun ahha

OpenStudy (jhannybean):

That is correct.

OpenStudy (anonymous):

Oh really? yay haha thanks so much!

OpenStudy (jhannybean):

\[5x-2x=2+15\]\[3x=17\]\[x=\frac{17}{3}\]